OpenStudy (anonymous):

I REALLY NEED HELP!!!! the drama club at Lincoln High School sells hot chocolate and coffee at the schools football games to make money for a special trip. at one game they sold $200 worth of hot drinks. they need to report how many of each type of drink they sold for their club records. Macha knows that they used 295 cups that night. if hot chocolate sells for 75 cents and coffee sells for 50 cents, how many of each type of hot drinks did they sell?

OpenStudy (anonymous):

it depends whether which was sold more 2 posibilities

OpenStudy (anonymous):

all i know is that i need help..

OpenStudy (anonymous):

im very confused....

OpenStudy (anonymous):

If number of hot choc cups is x and number of coffee cups is y: x + y = 295 .75x +.5y = 200 (I converted the cents to dollars) put this into your other equation: .75 (295 - y) + .5y = 200 221.25 - .75y + .5y = 200 221.25 - .25y = 200 21.25 = .25y y=21.25x4 y=85 So x = 295 - y x = 210

OpenStudy (anonymous):

oh thanx!!!

OpenStudy (anonymous):

hope it made sense....glad to help!